# RSEND: Retinex-based Squeeze and Excitation Network with Dark Region Detection for Efficient Low Light Image Enhancement

### v1

@@ Abstract
Images captured under low-light scenarios often suffer from low quality. Previous CNN-based deep learning methods often involve using Retinex theory. Nevertheless, most of them cannot perform well in more complicated datasets like LOL-v2 while consuming too much computational resources. Besides, some of these methods require sophisticated training at different stages, making the procedure even more time-consuming and tedious. In this paper, we propose a more accurate, concise, and one-stage Retinex theory based framework, RSEND. RSEND first divides the low-light image into the illumination map and reflectance map, then captures the important details in the illumination map and performs light enhancement. After this step, it refines the enhanced gray-scale image and does element-wise matrix multiplication with the reflectance map. By denoising the output it has from the previous step, it obtains the final result. In all the steps, RSEND utilizes Squeeze and Excitation network to better capture the details. Comprehensive quantitative and qualitative experiments show that our Efficient Retinex model significantly outperforms other CNN-based models, achieving a PSNR improvement ranging from 0.44 dB to 4.2 dB in different datasets and even outperforms transformer-based models in the LOL-v2-real dataset.
Keywords:
low light image enhancement Squeeze and excitation network compact model
1 Introduction
@@
Low-light image enhancement aims to improve the visibility and perceptual quality of images captured in underexposed environments. This field addresses challenges such as noise, color distortion, and loss of detail, which commonly affect low-light images.
@@
Figure 1: Comparison of our RSEND against previous CNN based state-of-the-art methods including SID [3], RetinexNet [21], EnGAN [12], DRBN [24], KinD [29] on three datasets. Our method shows a marked improvement, as indicated by the red bars and the associated increase annotated above them.
@@
A large number of approaches have been proposed for low-light image enhancement. Traditional image enhancement methods mainly include Retinex theory-based algorithm and histogram equalization. However, histogram equalization sometimes leads to over-amplification of noise in relatively dark areas of an image as well as a loss of detail in brighter sections. This technique applies a global adjustment to the image’s contrast, which may not be suitable for images where local contrast variations are important for detail visibility. As for Retinex theory-based methods, they assume an image could be decomposed into illumination and reflectance, while preserving the reflectance, by only enhancing the illumination can they get the final enhanced image. However, such methods sometimes produce results that appear unnatural due to over-enhancement or incorrect color restoration. Additionally, these models might struggle with very dark regions where information is minimal, potentially leading to artifacts or noise amplification.
@@
With the advancement of deep learning, Convolutional Neural Networks (CNNs) have become a pivotal technology for enhancing images captured in low-light conditions. Various types of CNN architectures have been explored, mainly they are divided into two categories, Generative Adversarial Networks (GANs) and models inspired by Retinex theory. For GANs, the adversarial process helps improve the quality of the enhancement, making images look more natural while suffering from training instability, leading to artifacts or unrealistic results, especially under complex lighting conditions. As for Retinex theory-based deep learning models, they typically decompose an image into reflectance and illumination components. The models estimate these components separately, enhancing the illumination part to improve image visibility while preserving the reflectance to maintain color fidelity and details. One drawback of such models is their reliance on accurate decomposition, which can be challenging in complex lighting conditions, and they always suffer from the multi-stage training pipeline.
@@
Another problem is that previous methods always consume a large number of parameters, and heavy computational complexity is unaffordable in certain situations. For example, privacy concerns are paramount when processing sensitive images on local devices. If we have a compact and efficient network, we can ensure that data processing occurs on the device itself, especially for mobile applications where memory and processing power are limited. And when deploying models in the cloud, reducing cost is a critical factor. Smaller models require less computational power, thereby saving more energy and reducing the financial burden associated with cloud resources.
@@
To fix the aforementioned problems, we propose a novel method, RSEND, to deal with low-light image enhancement tasks. First, we follow the Retinex theory methodology, decomposing the image into the illumination map and reflectance map. We add a dark region detection module so that the illumination map can further go through a multi-scale, separate pathway before enhancing to locate the features that need to be enhanced. Then we go through our U-shape [19] enhancer, and refine our image for better details. After the enhanced grayscale image does element-wise multiplication with the reflectance map, we add the original image to maintain similarity. And finally, we denoise the output for a more visually pleasing result. By utilizing Squeeze-and-Excitation Blocks [10] in all the steps mentioned, the network can recalibrate channel-wise feature responses by explicitly modeling inter-dependencies between channels, capturing more image details while not increasing substantial computational cost. We will open source this work to facilitate future research. Link to the source code can be found in the supplementary materials.
Our contributions are summarized as follows:
@@
  • We use a residual learning way in the reconstruction step to make sure our output is similar to the original low-light image, maintaining high SSIM.
@@
  • Our model outperforms all other CNN-based low-light image enhancement networks and even transformer-based models while costing much fewer computational resources.
2 Related Work
2.1 Traditional Methods
@@
Traditional methods for low-light image enhancement, such as histogram equalization [11] [4] and gamma correction [7], focus on globally adjusting image contrasts or brightness. While these methods are simple and fast, they often overlook local context and can lead to unrealistic effects or artifacts, such as over-enhancement, under-enhancement in certain areas, or amplified noise. These limitations stem from their global processing nature, which does not account for local variations in light distribution within an image. As a result, while effective for moderate adjustments, they may struggle with images having complex light conditions or requiring nuanced enhancements.
2.2 Deep learning Methods
@@
With the fast development of deep learning, CNNs [24], [3], [29], [6] have been extensively used in low light image enhancement. EnlightenGAN [12] utilizes unsupervised learning for low-light image enhancement, leveraging a global-local discriminator structure to ensure detailed enhancement and incorporates attention mechanisms to refine areas needing illumination adjustment, but a potential drawback is the challenge of maintaining naturalness and avoiding over-enhancement, especially in images with highly variable light conditions. ZeroDCE [8] tackles low-light image enhancement through a novel deep curve estimation approach that dynamically adjusts the light enhancement of images without needing paired datasets, which introduces a lightweight deep network to learn enhancement curves directly from data. However, like many approaches using unpaired datasets, its performance may depend on the diversity and quality of the training data, potentially limiting its adaptability to unseen low-light conditions. Retinex-based deep learning models [21] focus on separating the illumination and reflectance components of an image, allowing for the manipulation of illumination while preserving the natural appearance of the scene. Nevertheless, these networks always suffer from muti-stage training pipelines and face challenges including maintaining color fidelity and avoiding artifacts.
2.3 Squeeze-and-Excitation Network
@@
The Squeeze-and-Excitation Network (SENet) [10] introduces a mechanism to recalibrate channel-wise feature responses adaptively by explicitly modeling interdependencies between channels. It squeezes global spatial information into a channel descriptor using global average pooling, then captures channel-wise dependencies through a simple gating mechanism, and finally excites the original feature map by reweighting the channels. The SENet approach has been applied to a wide range of tasks beyond low-light image enhancement, such as image classification, object detection, semantic segmentation, and medical image analysis. It has also been used in this area, The Multiscale Low-Light Image Enhancement Network with Illumination Constraint [6] utilizes SENet for low-light image enhancement. However, it consumes substantial computational resources for training and inference while not reaching a desired PSNR value.
@@
3 RSEND: Efficient Low-Light Image Enhancement
In this section, we introduce our method RSEND, as shown in Fig. 2. We then provide the details of our work.
3.1 End-to-end Retinex-based Model
As we mentioned previously, Retinex theory has been extensively used in this area, A low light image $S\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $I\in\mathbb{R}^{H\times W}$
@@ 3.1 End-to-end Retinex-based Model
the $\circ$ here represents element-wise multiplication. Similar to the previous approaches, we decompose the image into reflectance and illumination that is designed to enhance. Unlike RetinexNet [21] to manually separate the first three channels as $R$ and last channel as $I$ after a few convolutional layers, we are inspired by Retinexformer [2] and like many other models did, after feature extraction, use two different layers in the end, one outputs three channels and the other outputs one channel, to separate $R$ and $I$ and use SEBlock in the middle for better feature extraction.
3.2 The Illumination Procedure
Instead of directly enhancing the illumination map, we add a few more illumination constraints. Before enhancing, we design a dark region detection module, using convolutions at Illumination map with kernel 3x3 and 5x5 and stride set as 2 to capture features at different scales, then apply sigmoid activation to generate attention maps, which weights the importance of each region needs to be enhanced. By upsampling the features from different scales to the original size and doing concatenation with the original feature map, we come up with a feature map that has channels*3 depth and features from both attention-augmented pathways. In the end, we apply a 1x1 kernel size convolutional layer to the concatenated multi-scale features to reduce the channel dimensions to $I\in\mathbb{R}^{H\times W\times 1}$. By applying this module, the network is expected to pay more attention to the darker areas that need enhancement.
@@ 3.2 The Illumination Procedure
3.3 Reconstruction and Denoising Phase
@@
3.4 Compact Network
For low-light image enhancement tasks, previous CNN-based models always add layers and increase depth for more feature representation. However, this usually does not yield promising results while largely increasing computational costs. Our RSEND framework exemplifies the principle of reducing computational costs by integrating key design choices that promote efficiency. Firstly, the use of Squeeze-and-Excitation blocks allows the model to perform dynamic channel-wise feature recalibration, which significantly boosts the representational power of the network without a proportional increase in parameters. Secondly, we carefully design the depth of the network so that each layer contributes meaningfully to the feature extraction process. Previous works always reach a depth of 512 in the bottleneck of the enhancer, while in our work, the bottleneck only has 64 channels and in other modules of the network the depth is restrained to 32. The result in Table 1 shows that it is possible to build powerful yet compact models with a fraction of the parameters in the realm of low-light image enhancement.
3.5 Loss Function
Our training objective consists of several loss components designed to capture different aspects of image quality in low-light image enhancement tasks. Each term contributes to a specific characteristic that we aim to imbue in the enhanced output.
3.5.1 Spatial Consistency Loss
The Spatial Consistency Loss $\mathcal{L}_{spa}$ enhances the spatial coherence of the enhanced image by maintaining the difference of neighboring regions between the input and its enhanced version:
$\mathcal{L}_{spa}=\frac{1}{K}\sum_{i=1}^{K}\sum_{j\in\Omega(i)}\left(\left|Y_{i}-Y_{j}\right|-\left|I_{i}-I_{j}\right|\right)^{2},$ (6)
where $K$ represents the number of local regions, $\Omega(i)$ denotes the four neighboring regions (top, down, left, right) centered at region $i$. $Y$ and $I$ are the average intensity values of the local region in the enhanced and input images, respectively.
3.5.2 Exposure Control Loss
The Exposure Control Loss $\mathcal{L}_{exp}$ aims to correct under-/over-exposed regions by controlling the exposure level. It measures the distance between the average intensity of a local region and a predefined well-exposedness level $E$:
$\mathcal{L}_{exp}=\frac{1}{M}\sum_{k=1}^{M}\left|Y_{k}-E\right|,$ (7)
where $M$ is the number of non-overlapping local regions, $Y_{k}$ is the average intensity value of the $k^{th}$ local region in the enhanced image, and $E$ is the target exposure level. In our case, we set $E$ as 0.6.
3.5.3 Color Constancy Loss
The Color Constancy Loss $\mathcal{L}_{col}$ corrects potential color deviations by adhering to the Gray-World hypothesis [1], which assumes that the average color in a scene is achromatic. This loss ensures that the average color of the sensor channels converges towards gray and prevents potential color deviation over the entire image:
$\mathcal{L}_{col}=\sum_{\varepsilon\in\{(R,G),(R,B),(G,B)\}}\left(J_{p}^{\varepsilon}-J_{q}^{\varepsilon}\right)^{2},$ (8)
where term $J^{\varepsilon}$ represents the average intensity of a particular pair of color channels within the image. The indices $p$ and $q$ here refer to different spatial regions of the image where color constancy is being compared, and the summation runs over all combinations of the RGB channels.
3.5.4 Perceptual Loss ($\mathcal{L}_{vgg}$) [13]
To ensure that the enhanced image is perceptually similar to the well-lit ground truth, we employ a perceptual loss using the feature maps from a pre-trained VGG-19 network:

### v2

@@ Abstract
Images captured under low-light scenarios often suffer from low quality. Efficient low-light image enhancement with mobile computing has become an urgent need. Previous CNN-based low-light image enhancement methods often involve using Retinex theory. Nevertheless, most of them do not perform well in complicated datasets like LOL-v2 while using too much computational resources. Besides, some of these methods require sophisticated training at different stages, making the procedure even more time-consuming and tedious. In this paper, we propose an accurate, concise, and one-stage Retinex theory-based framework with a novel dark region detection module and Squeeze and Excitation blocks for enhanced detail retention, RSEND, for efficient low-light image enhancement. RSEND first divides the low-light image into the illumination map and reflectance map, then detects the different dark regions in the illumination map and performs light enhancement. After this step, it refines the enhanced gray-scale image and does element-wise matrix multiplication with the reflectance map. By denoising the output it has from the previous step, it obtains the final result. In all the steps, RSEND utilizes Squeeze and Excitation network to better capture the details. Comprehensive quantitative and qualitative experiments show that our efficient Retinex model significantly outperforms other CNN-based state-of-the-art models, achieving a PSNR improvement ranging from 1.69 dB to 3.63 dB in different datasets. Compared to Transformer-based models, RSEND achieves higher PSNR values ranging from 1.22 dB to 2.44 dB in the LOL-v2-real dataset. Importantly, RSEND achieves these performance improvements with remarkable efficiency, utilizing only 0.41 million parameters, which represents a substantial reduction (3.93–9.78$\times$) in computational resources compared to existing state-of-the-art methods. The code can be found at https://github.com/jeffconqueror/RSEND/tree/main.
I Introduction
@@
Low-light image enhancement aims to improve the visibility and perceptual quality of underexposed images captured in underlit environments. This enhancement problem is challenging as it requires careful handling of noise, color distortion, loss of details, etc. while keeping the model training and inference computing footprint reasonably low.
@@
Figure 1: Comparison of our RSEND against previous CNN based state-of-the-art methods including SID [1], RetinexNet (RTXNet) [2], EnGAN [3], DRBN [4], KinD [5] on three datasets: LOL-v1, LOL-v2-real, and LOL-v2-syn. Our RSEND flow achieves 1.69 dB to 3.63 dB improvements over the best previous works in terms of PSNR, as indicated by the orange +dB annotations.
@@
Many prior approaches have been proposed for low-light image enhancement in literature. Previous image enhancement methods mainly include Retinex theory-based [6] [2] algorithm and histogram equalization [7]. However, both techniques have their own drawbacks. Histogram equalization sometimes leads to over-amplification of noise in relatively dark areas of an image as well as a loss of detail in brighter sections [8]. This technique applies a global adjustment to the image’s contrast, which may not be suitable for images where local contrast variations are important for detail visibility. As for Retinex theory-based methods, they assume an image could be decomposed into illumination and reflectance, while preserving the reflectance, by only enhancing the illumination can they get the final enhanced image. However, such methods sometimes produce results that appear unnatural due to over-enhancement or incorrect color restoration. Additionally, these models might struggle with very dark regions where information is minimal, potentially leading to artifacts or noise amplification.
@@
With the advancement of deep learning, convolutional neural networks (CNNs) have become a pivotal technology for enhancing images captured in low-light conditions. Various types of CNN-based architectures have been explored. They are mainly divided into two categories, generative adversarial networks (GANs) [3] and models inspired by Retinex theory [2] [9] [8]. For GANs, the adversarial process helps improve the quality of the enhancement, making images look more natural. However, GAN-based methods suffer from training instability, leading to artifacts or unrealistic results, especially under complex lighting conditions [10]. As for Retinex theory-based deep learning models, they typically decompose an image into reflectance and illumination components. The models estimate these components separately, enhancing the illumination part to improve image visibility while preserving the reflectance to maintain color fidelity and details. One drawback of such models is their reliance on accurate decomposition, which can be challenging in complex lighting conditions, and they always suffer from the multi-stage training pipeline.
@@
Another problem is that previous methods typically require building models with a large number of parameters, which leads to heavy computational complexity that is unaffordable in certain situations, e.g., image enhancement on mobile devices. Besides computational considerations, privacy concerns are paramount when processing sensitive images on local devices. If we have a compact and efficient network, we can ensure that image processing occurs locally on the device itself, especially for mobile applications where memory and processing power are limited. Not only for mobile and edge devices, reducing computing cost is also a critical need for deploying models in the cloud. Smaller models require less computational power, thereby saving more energy and reducing the financial burden associated with cloud resources.
@@
To fix the aforementioned problems, we propose a novel method, RSEND, for efficient low-light image enhancement with high quality and low computing cost. First, RSEND adopts the Retinex theory methodology and decomposes the low-light image into the illumination map and the reflectance map. Prior works didn’t pay attention to locating the areas that need the most enhancement. We solve this by adding a dark region detection module so that the illumination map can further go through a multi-scale, separate pathway to locate the features that require enhancing, before the actual enhancement. Then, RSEND goes through our custom U-shape [11] enhancer, and refine the input image for better details. After the enhanced grayscale image does element-wise multiplication with the reflectance map, we add the original image back to maintain similarity. And finally, RSEND denoises the output for a more visually pleasing result. By utilizing Squeeze-and-Excitation Blocks [12] in all the steps mentioned, the network can recalibrate channel-wise feature responses by explicitly modeling inter-dependencies between channels, capturing more image details while not increasing substantial computational cost. Fig. 1 shows the peak signal-to-noise ratio (PSNR) result of our RSEND flow compared against state-of-the-art works on three representative datasets. Our RSEND flow achieves 1.69 dB to 3.63 dB improvements over the best previous works. We will open source this work to facilitate future research. Link to the source code can be found in the supplementary materials.
The major contributions of this work can be summarized as follows:
@@
  • We use a residual learning way in the reconstruction step to make sure our output is similar to the original low-light image, maintaining high structural similarity index measure (SSIM).
@@
  • Our RSEND model outperforms all other CNN-based low-light image enhancement networks by up to 3.63 dB and even Transformer-based models by up to 2.44 dB while utilizing only 0.41 million parameters, which is 3.93 – 9.78$\times$ reduction in model size.
II Related Works
II-A Traditional Methods
@@
Traditional methods for low-light image enhancement, such as histogram equalization [7] [13] and gamma correction [14], focus on globally adjusting image contrasts or brightness. While these methods are simple and fast, they often overlook local context and can lead to unrealistic effects or artifacts, such as over-enhancement or under-enhancement in certain areas, or amplified noise. These limitations stem from their global processing nature, which does not account for local variations in light distribution within an image. As a result, while effective for moderate adjustments, they may struggle with images having complex light conditions or requiring nuanced enhancements.
II-B Deep Learning Methods
@@
With the fast development of deep learning, CNNs [4, 1, 5, 15] have been extensively used in low light image enhancement. EnlightenGAN [3] utilizes unsupervised learning for low-light image enhancement, leveraging a global-local discriminator structure to ensure detailed enhancement and incorporates attention mechanisms to refine areas needing illumination adjustment, but a potential drawback is the challenge of maintaining naturalness and avoiding over-enhancement, especially in images with highly variable light conditions. ZeroDCE [16] tackles low-light image enhancement through a novel deep curve estimation approach that dynamically adjusts the light enhancement of images without needing paired datasets, which introduces a lightweight deep network to learn enhancement curves directly from data. However, like many approaches using unpaired datasets, its performance may depend on the diversity and quality of the training data, potentially limiting its adaptability to unseen low-light conditions. Retinex-based deep learning models [2] focus on separating the illumination and reflectance components of an image, allowing for the manipulation of illumination while preserving the natural appearance of the scene. Nevertheless, these networks always suffer from muti-stage training pipelines and face challenges including maintaining color fidelity and avoiding artifacts.
II-C Squeeze-and-Excitation Network
@@
The Squeeze-and-Excitation Network (SENet) [12] introduces a mechanism to recalibrate channel-wise feature responses adaptively by explicitly modeling interdependencies between channels. It squeezes global spatial information into a channel descriptor using global average pooling, then captures channel-wise dependencies through a simple gating mechanism, and finally excites the original feature map by reweighting the channels. The SENet approach has been applied to a wide range of tasks beyond low-light image enhancement, such as image classification, object detection, semantic segmentation, and medical image analysis. The MLLEN-IC work (Multiscale Low-Light Image Enhancement Network with Illumination Constraint) [15] utilizes SENet for low-light image enhancement. The paper presents a comprehensive solution by combining a multiscale network architecture with an illumination constraint, and by using SENet, the model better restores the color and details of the image. However, it consumes substantial amount of computational resources for training and inference while only has an average PSNR value of 15.11 and SSIM value of 0.56.
@@
III RSEND: Efficient Low-Light Image Enhancement
In this section, we introduce our proposed low-light image enhancement framework, RSEND. The overall architecture of RSEND is presented in Fig. 2.
III-A End-to-end Retinex-based Model
As we mentioned previously, RSEND first applies Retinex theory [6] to the input low-light image. A low light image $S\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $I\in\mathbb{R}^{H\times W}$
@@ III-A End-to-end Retinex-based Model
where the $\circ$ operator represents element-wise multiplication along the $H\times W$ dimensions and repeated across RGB channels. Similar to the previous approaches, we decompose the image into reflectance and illumination in the first step. Although many current mainstream methods have similar decomposition step, we still find there is room to improve the quality in both reflectance and illumination steps. Unlike RetinexNet [2] which manually separates the first three channels as $R$ and last channel as $I$ after a few convolutional layers, after feature extraction, we use two different layers in the end, one outputs three channels and the other outputs one channel, to separate $R$ and $I$ and use SEBlock [12] in the middle for better feature extraction.
III-B Dark Region Detection Module
Instead of directly enhancing the illumination map, we introduce a novel dark region detection module designed to emphasize areas requiring greater enhancement. As shown in Fig. 3, the left side presents the illumination map before dark region detection, where uniform enhancement results in limited visibility improvement. The right side illustrates the effect after applying our dark region detection module, highlighting enhanced regions more effectively.
We use convolutions at Illumination map with kernel 3x3 and 5x5 and stride set as 2 to capture features at different scales, then apply sigmoid activation to generate attention maps, which weights the importance of each region needs to be enhanced. By upsampling the features from different scales to the original size and doing concatenation with the original feature map, we come up with a feature map that has channels*3 depth and features from both attention-augmented pathways. In the end, we apply a 1x1 kernel size convolutional layer to the concatenated multi-scale features to reduce the channel dimensions to $I\in\mathbb{R}^{H\times W\times 1}$. By applying this module, the network is expected to pay more attention to the darker areas that need enhancement.
@@ III-B Dark Region Detection Module
Before Dark Region Detection
After Dark Region Detection
Figure 3: Effect of Dark Region Detection. The left image shows the illumination map before dark region detection, while the right image demonstrates enhanced focus on darker areas after applying the module.
III-C Illumination Optimization
@@
III-D Reconstruction and Denoising Phase
@@
III-E Compact Network
For low-light image enhancement tasks, previous CNN-based models always add layers and increase depth for more feature representation. However, this usually does not yield promising results while largely increasing computational costs. Our RSEND framework exemplifies the principle of reducing computational costs by integrating key design choices that promote efficiency. Firstly, the use of Squeeze-and-Excitation blocks allows the model to perform dynamic channel-wise feature recalibration, which significantly boosts the representational power of the network without a proportional increase in parameters. Secondly, we carefully design the depth of the network so that each layer contributes meaningfully to the feature extraction process. Previous works always reach a depth of 512 in the bottleneck of the enhancer, while in our work, the bottleneck only has 64 channels and in other modules of the network the depth is restrained to 32. The result in Table I shows that it is possible to build powerful yet compact models with a fraction of the parameters in the realm of low-light image enhancement.
IV Experiment
IV-A Datasets and Implementation details
We evaluate our model on four paired datasets. The LOL-v1, LOL-v2-real captured, LOL-v2-synthetic, and SID datasets, which training and testing are split into 485:15, 689:100, 900:100, and 2564:133. We resize the training image to 224x224, and implement our framework with PyTorch on two NVIDIA 4090 GPUs with a batch size of 8. The model is trained using the AdamW optimizer with initial hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Training is conducted for 750 epochs, where the learning rate is initially set to $1\times 10^{-8}$ and increases to $2\times 10^{-5}$ over the first 75 epochs during a warmup phase. Subsequently, the learning rate is maintained at $2\times 10^{-5}$ until the 600th epoch, after which it follows a cosine annealing schedule down to $1\times 10^{-8}$ towards the end of the training at 750 epochs. To ensure that the enhanced image is perceptually similar to the well-lit ground truth, we employ a perceptual loss [21] using the feature maps from a pre-trained VGG-19 network, and we adopt the peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) as the evaluation metrics.